OpenStudy (anonymous):

What is the equation of the line perpendicular to 2x – 3y = 13 that passes through the point (–6, 5)?

OpenStudy (anonymous):

just find the midpoint and slope

OpenStudy (anonymous):

I'm not sure how, lol.

OpenStudy (pinksapphire):

Solve the equation by getting y by itself first.

OpenStudy (pinksapphire):

2x-3y=13 -2x -2x -3y=12-2x -3y/-3=12-2x/-3 y=-4+2/3x Slope: 2/3x y-intercept: -4

OpenStudy (anonymous):

So, will the answer be y=2/3x-1? Sorry, I'm terrible at math, lol.

OpenStudy (pinksapphire):

No, the equation you wrote is parallel to the other one.

OpenStudy (anonymous):

Oh, right, I'm looking for perpendicular. Lol. y=2/3x+9 then, right? I think I get it now.

OpenStudy (pinksapphire):

No, that's parallel too.

OpenStudy (pinksapphire):

Maybe this will help: http://www.mathsisfun.com/algebra/line-parallel-perpendicular.html

OpenStudy (anonymous):

Thank you for the link, I think it'll be really helpful. And thanks for the help!

OpenStudy (pinksapphire):

Okay :)